ABS stock is selling for $30 a share. Price of a 6-month call option with a strike price of $32 is $0.50. Risk-free rate is 0.5% per month. What is the price of a 6-month put option with a strike price of $32?
Answer #1

Using Put Call Parity Equation,

C + X/(1 + r)t = S + P

0.50 + 32/(1.005)0.50 = 30 + P

P = $2.42

Price of Put Option = $2.42